OpenStudy (mendicant_bias):

Alright, got another Electrochemical cell problem. Posted below momentarily.

OpenStudy (mendicant_bias):

"How many grams of chromium metal are plated out when a constant current of 8.00 A is passed through an aqueous solution containing Cr3+ ions for 40.0 minutes?"

OpenStudy (mendicant_bias):

@JFraser , I'm not really sure where to go with this one.

OpenStudy (mendicant_bias):

I know I'm going to have to do some dimensional analysis/unit conversion stuff, but I'm just sort of eyeballing here.

OpenStudy (mendicant_bias):

\[40.0 \rm \ minutes = (60)(40)=2,400\]

OpenStudy (mendicant_bias):

(seconds)

OpenStudy (jfraser):

this is a rather long dimensional analysis question, you're right but you need to use the UNITS to tell you where the numbers go, not the other way around

OpenStudy (jfraser):

put the units IN the problem, like this:

OpenStudy (jfraser):

\[40 \cancel {minutes} * \frac{60 seconds}{1\cancel{minute}}\]

OpenStudy (mendicant_bias):

\[2,400 \rm \ s = t\]

OpenStudy (mendicant_bias):

\[8.0 \rm A = 8.0 \frac{C}{s}; \]\[\rm \frac{8.0 \ C }{s}\cdot 2,400 \ s = 19,200 \ C\]

OpenStudy (jfraser):

You can keep on going this way, doing the math one step at a time, OR:

OpenStudy (jfraser):

\[40 minutes * \frac{60seconds}{1 minute}*\frac{8.00Coulomb}{1 second}*\frac{1 mol \space e^{-1}}{96,500Coulombs}*\frac{1mol \space Cr}{3mol \space e^{-1}}*\frac{51.99g \space Cr}{1mol \space Cr}\]

OpenStudy (mendicant_bias):

Your LaTeX got cut off :/

OpenStudy (jfraser):

damn

OpenStudy (jfraser):

the last two are \[\frac{1mol \space Cr}{3mol \space e^{-1}}*\frac{51.99g \space Cr}{1mol \space Cr}\]

OpenStudy (mendicant_bias):

Alright, but I think I got the gist of it conceptually, turn quantity given into units of charge, divide by Faraday's constant, and divide by the number of electrons associated with the half reaction.....and then multiply by its molar mass

OpenStudy (jfraser):

you're replacing one kind of unit with another kind of unit, just using the data they give you. the units tell you everything you need, even without the numbers

OpenStudy (mendicant_bias):

Yeah, makes sense. Alright. I'm going to move on to another problem, thank you.